OpenStudy (anonymous):

At Heard High School, the probability that a student takes Advanced Weight Training and Advanced Spanish is 0.099. The probability that a student takes Advanced Weight Training is 0.71. What is the probability that a student is taking Advanced Spanish if the same student is taking Advanced Weight Training?

OpenStudy (hoblos):

use p(B/A) = P(AnB)/P(A)

OpenStudy (hoblos):

so the answer would be 0.099/0.71 = 0.139